VIB–ROTATIONAL ANALYSIS ON 2ν_{1} + ν_{5}, ν_{1} + ν_{2} + 3ν_{5}, and 2ν_{2} + 5ν_{5} COMBINATION BANDS OF ACETYLENE USING A GaAs SEMICONDUCTOR LASER

\underline{M.SUZUKI}, T.YAMASHITA, T.OHTSUKA, T.YUKIYA, AND N.NISHIMIYA, Dept. of Electronic Engineering, Tokyo Institute of Polytechnics, Iiyama 1583, Atsugi City, 243-0297 Kanagawa, Japan.

The vib–rotational spectrum of the 2ν_{1}+ν_{5}^{l=1}, ν_{1}+ν_{2}+3ν_{5}^{l=1}, and 2ν_{2}+5ν_{5}^{l=1} combination bands of acetylene was measured using a GaAs semiconductor laser (New Focus 6148) at the wavenumber region from 7400 to 7600 cm^{−1} and the P–, R– and Q–branches were assigned. The wavenumber measuring system with a wavelength meter (Anritsu MF9630A) and a confocal interferometer of 15 cm was calibrated against the H_{2}O lines.^{1} The formulation describing the spectral line positions by a third order polynomial of J(J+1) with the effective spectroscopic parameters of ν_{0}^{eff}, B_{v}^{eff}, D_{v}^{eff}, and H_{v}^{eff} was obtained using second order perturbation theory to take the l–type resonance effect into consideration. The spectroscopic constants and the l–type doubling constants were calculated from the effective rotational parameters determined using a least squares fitting procedure for the e– and f–levels by the aid of the molecular constants reported by ref. 2 and 3.

1. R.A.Toth, Appl.Opt., 33, 4851 (1994).

2. M.Herman, T.R.Huet, Y.Kabbadj, and J.V.Auwera, Mol.Phys., 72, 75 (1991).

3. J.J.Hillman, D.E.Jennings,G.W.Halsey,S.Nadler,and E.W.Blass, J.Mol.Spectrosc., 146, 389(1991).